# Activation of entanglement in generalized entanglement swapping

### v1

@@ III CONCLUSIONS
In this study, we explore entanglement activation in a generalized entanglement swapping process involving two Bell pairs and generalized measurements. The conventional wisdom suggests that entangled measurements are both necessary and sufficient for establishing entanglement between distant parties. Previous research, as noted in references terhal ; vincenzo ; yokoyama , has established the necessity and sufficiency of entangled measurements to obtain an entangled state between Alice and Charlie when the initial states are Bell pairs, and measurements are generalized measurements, known as POVM.
@@
In conclusion, expanding the protocol to include non-maximally entangled states would enhance its applicability, while extending it to cover a broader range of scenarios holds promise for further advancements in the field. Future research efforts should prioritize refining entanglement activation protocols, exploring innovative measurement strategies, and broadening the protocol’s applicability to diverse scenarios, including nonlocality activation klobus and network nonlocality bran1 ; bran2 . By addressing these research directions, we can deepen our understanding of entanglement generation and distribution, leading to advancements in quantum communication and computation.
@@ Acknowledgements
The authors express gratitude to Prof. Somshubhro Bandyopadhyay for invaluable discussions. PB acknowledges Saronath Halder for their insightful and stimulating discussions.
@@
Appendix A Inseparable quantum operation condition of an arbitrary two qubits measurement:
@@
$=$ $\frac{1}{\mbox{Tr}(\Pi_{n})}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j}a^{\ast}_{n\beta|k^{\prime}l}|ik\rangle\langle i^{\prime}k^{\prime}|_{12}$ (77)
@@
$\rho_{12|n}^{2}=\frac{1}{\mbox{Tr}(\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime},p,q,\\ p^{\prime},q^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|pj}a_{n\eta|p^{\prime}j}a^{*}_{n\eta|q^{\prime}l}a_{n\gamma|ql}|ik\rangle\langle i^{\prime}k^{\prime}|pq\rangle\langle p^{\prime}q^{\prime}|$ (83)
@@
$\rho_{12|n}^{2}=\frac{1}{\mbox{Tr}(\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime},\\ p^{\prime},q^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|p^{\prime}j}a^{*}_{n\eta|q^{\prime}l}a_{n\gamma|k^{\prime}l}|ik\rangle\langle p^{\prime}q^{\prime}|$ (89)
@@
$\mbox{Tr}(\rho_{12|n}^{2})=\frac{1}{\mbox{Tr}(\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}.$ (94)
@@
The entanglement in the $12|34$ bipartition of the post-measurement state when the $\Pi_{n}$ element of any arbitrary two-qubits measurement (Eq.71) is performed on the (2,3) pair can be represented using Eq.94
$\mathcal{C}_{n|12vs34}(|\Phi_{n})\rangle\langle\Phi_{n}|)=\sqrt{\frac{4}{3}\left[1-\frac{1}{\mbox{Tr}(\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}\right]}$ (99)
Appendix B Negativity of the post-measurement state after sequential measurements
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta|i_{8}j_{4}}|i_{4}j_{3}\rangle\langle i_{8}j_{7}|$ (114)
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{3}}a^{*}_{n\eta|i_{4}j_{4}}$ (122)
@@
$\times$ $*\displaystyle\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}}$ (131)
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}},$ (139)
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}},$ (147)
@@
$\times$ $*.\displaystyle\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}}.$ (155)
So, the negativity of the final post-measurement state $\rho_{14|nm}$ (Eq.29) can written as:
@@
The negativity of the post-measurement state of the (1,4) pair can be calculated by using Eq.121 for $X$ and Eq.123 for $Y$.

### v2

@@ III CONCLUSIONS
In this study, we explore entanglement activation in a generalized entanglement swapping process involving two Bell pairs and generalized measurements. Conventional wisdom suggests that entangled measurements are required to establish entanglement between distant parties.
@@
In conclusion, expanding the protocol to include non-maximally entangled states would enhance its applicability, while extending it to cover a broader range of scenarios holds promise for further advancements in the field. We have additionally addressed this matter by considering pure non-maximally entangled states ($\cos{\theta}|00\rangle+\sin{\theta}|11\rangle$) as initial states for the qubit pairs (1,2) and (3,4). This state exhibits entanglement for $0<\theta<\frac{\pi}{2}$, reaching maximum entanglement at $\theta=\frac{\pi}{4}$ with a concurrence of $2\cos\theta\sin\theta$. We have then performed sequential measurements using Bell basis with white noise as the first POVM measurement Eq.II.1.2, followed by a second measurement Eq.II.1.2. After the first measurement, the entanglement of the output state in the (1,4) pair depends on both $\theta$ and $\lambda$. However, we did not observe a state that remains entangled across the entire range of $\lambda$ for a fixed value of $\theta$. Conversely, the sequential application of these two measurements results in the (1,4) pair being entangled across the entire range of $\lambda$ for $0<\theta<\frac{\pi}{2}$. Therefore, our protocol is also applicable in scenarios involving pure non-maximally entangled states for entanglement distribution. Despite the challenges in obtaining a compact expression for the post-measurement state in (1,4) pair under arbitrary measurements with non-maximally entangled initial states, our findings suggest promising avenues for future research.
Future research efforts should prioritize refining entanglement activation protocols, exploring innovative measurement strategies, and broadening the protocol’s applicability to diverse scenarios, including nonlocality activation klobus and network nonlocality bran1 ; bran2 . By addressing these research directions, we can deepen our understanding of entanglement generation and distribution, leading to advancements in quantum communication and computation.
@@ Acknowledgements
The authors express gratitude to Prof. Somshubhro Bandyopadhyay for fruitful discussions. PB acknowledges Saronath Halder for helpful discussions.
@@
Appendix A Post-measurement state of (1,2) and (3,4) pairs
The density of state of the above Eq.31 is
$\rho_{1234|n}$ $=$ $|\Phi_{n}\rangle\langle\Phi_{n}|$ (32)
From the orthogonality condition of the orthonormal basis, we get
$=$ $\delta_{\beta\alpha}$
which reduce to
$\sum_{i,j=0}^{1}a^{*}_{n\beta|ij}a_{n\alpha|ij}=\delta_{\beta\alpha}$ (34)
$\rho_{12|n}$ $=$ $\frac{1}{4p_{n}}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j}a^{\ast}_{n\beta|k^{\prime}l}|ik\rangle\langle i^{\prime}k^{\prime}|_{12}$
$=$ $\frac{1}{\mbox{Tr}(\Pi_{n})}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j}a^{\ast}_{n\beta|k^{\prime}l}|ik\rangle\langle i^{\prime}k^{\prime}|_{12}$
From the above Eq.A, and Eq.36 one can easily compute the post-measurement states of $n$-th outcome between different pairs (1,2), and (3,4) respectively after the $\mathcal{M}_{1}$ POVM measurement.
Appendix B Inseparable quantum operation condition of an arbitrary two qubits measurement:
@@
$=$ $\frac{1}{\mbox{Tr}(\Pi_{n})}\sum_{\alpha,\beta=1}^{4}\sum_{\begin{subarray}{c}i,j,k,l,\\ i^{\prime},k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}a^{*}_{n\alpha|ij}a_{n\alpha|kl}a_{n\beta|i^{\prime}j}a^{\ast}_{n\beta|k^{\prime}l}|ik\rangle\langle i^{\prime}k^{\prime}|_{12}$
@@
$\rho_{12|n}^{2}=\frac{1}{(\mbox{Tr}\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime},p,q,\\ p^{\prime},q^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|pj}a_{n\eta|p^{\prime}j}a^{*}_{n\eta|q^{\prime}l}a_{n\gamma|ql}|ik\rangle\langle i^{\prime}k^{\prime}|pq\rangle\langle p^{\prime}q^{\prime}|$ (39)
@@
$\rho_{12|n}^{2}=\frac{1}{(\mbox{Tr}\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime},\\ p^{\prime},q^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|p^{\prime}j}a^{*}_{n\eta|q^{\prime}l}a_{n\gamma|k^{\prime}l}|ik\rangle\langle p^{\prime}q^{\prime}|$ (40)
@@
$\mbox{Tr}(\rho_{12|n}^{2})=\frac{1}{(\mbox{Tr}\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}.$ (41)
@@
The entanglement in the $12|34$ bipartition of the post-measurement state when the $\Pi_{n}$ element of any arbitrary two-qubits measurement is performed on the (2,3) pair can be represented using Eq.41
$\mathcal{C}_{n|12vs34}(|\Phi_{n})\rangle\langle\Phi_{n}|)=\sqrt{\frac{4}{3}\left[1-\frac{1}{(\mbox{Tr}\Pi_{n})^{2}}\sum_{\begin{subarray}{c}\alpha,\beta,\\ \gamma,\eta=1\end{subarray}}^{4}\sum_{\begin{subarray}{c}i,j,i^{\prime},k,\\ l,k^{\prime}=0\end{subarray}}^{1}\sqrt{\pi_{n\alpha}}\sqrt{\pi_{n\beta}}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\alpha|ij}a_{n\beta|i^{\prime}j}a^{*}_{n\beta|k^{\prime}l}a_{n\alpha|kl}a^{*}_{n\gamma|i^{\prime}j}a_{n\eta|ij}a^{*}_{n\eta|kl}a_{n\gamma|k^{\prime}l}\right]}$ (42)
Appendix C Negativity of the post-measurement state after sequential measurements
Negativity of a bipartite state $\rho_{AB}$ is defined as vidal
$\mathcal{N}(\rho_{AB})=\frac{\|\rho_{AB}^{T_{B}}\|-1}{2}$ (43)
@@ Appendix C Negativity of the post-measurement state after sequential measurements
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta|i_{8}j_{4}}|i_{4}j_{3}\rangle\langle i_{8}j_{7}|$
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{3}}a^{*}_{n\eta|i_{4}j_{4}}$
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}}$
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}},$
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}},$
@@
$\times$ $\mu_{m\delta}\sqrt{\pi_{n\gamma}}\sqrt{\pi_{n\eta}}a^{*}_{n\gamma|i_{5}j_{5}}b_{m\delta|k_{3}l_{3}}b^{*}_{m\delta|k_{4}l_{4}}a_{n\eta|i_{6}j_{6}}a_{n\beta|i_{3}j_{7}}a^{*}_{n\eta||i_{8}j_{4}}.$
@@
where $X$ and $Y$ are given above in Eq.47 and Eq.48 respectively. We get a relation between the negativity of the post-measurement state and the measurement parameters of two successive measurements.